Question 1 3 pts An acquirer with a P/E ratio of 13 and earnings of $1.95 seeks to take over another target firm with value $17.6 and P/E ratio 16. What is the new merged firms P/E? Question 2 3 pts Your firm is expected to earn $1.63/share next year and has a cost of capital of 17.4%. Assume these earnings resemble a perpetuity with growth rate 6.9%. what is its price/earnings ratio? DQuestion 3 3 pts What is an important problem with the equity-price-to-sales ratio? It can be negative It cannot be large It may reflect future sales growth It cannot be small It decrease one-to-one with leverage Please answer them Carry out calculations to at least 4 decimal places. Enter percentages as whole numbers. Example: 3.03% should be entered as 3.03. Do not include commas or dollar signs in numerical answers.
Answer #1

I have answered only question 2 . Hope this helps

2) expected earning next year = D1 = 1.63/share

growth rate = g = 6.9%

cost of capital = r = 17.4%

price = D1/(r-g) = 1.63/(17.4%-6.9%) = 15.52

price to earnings = price/earnings per share = 15.52/1.63 = 9.52